# ADAPTIVE-$\lambda$ SISS in MACHINE UNLEARNING for DDPM & VAE

### v1

@@ Abstract
Machine Unlearning (MU) is critical for large generative models, such as VAEs[1] and DDPMs[4], to enforce the right to be forgotten and ensure safe content generation without expensive retraining. Existing methods, like the Static-$\lambda$ SISS framework[6] for Diffusion Models, rely on a fixed data mixing parameter $\lambda$, which is suboptimal as the necessary unlearning rate varies dynamically across samples and training stages.
We introduce ADAPTIVE-$\lambda$ SISS, a novel approach that transforms the mixing parameter $\lambda$ from a fixed constant into a latent variable whose optimal value is dynamically inferred at each unlearning step. We leverage a Variational Inference-based objective[2], employing a small neural network to learn an adaptive posterior distribution $q_{\phi}(\lambda)$. This network conditions on context features derived from the SISS loss components, enabling simultaneous and robust optimization of the diffusion model and the $\lambda$-inference mechanism.
We also explore the application of this adaptive concept to score-function-based unlearning, proposing a multi-class extension for the Score Forgetting Distillation (SFD) framework[7]. Empirical results on an augmented MNIST dataset demonstrate that ADAPTIVE-$\lambda$ SISS significantly improves over the original SISS, yielding superior removal of forgotten content while better maintaining the generative quality on the retain set.
@@ 1 Introduction
Practical motivations for unlearning span several dimensions. Individuals may demand that facial images resembling them in datasets such as CelebA be removed from the learned model. More broadly, erasing semantic concepts or class labels enables a pathway toward controllable editing and concept-level modification within generative models. In addition, unlearning supports safe guidance, where a model avoids sampling from inappropriate or undesirable regions of the data distribution. These considerations, together with the infeasibility of retraining large models from scratch, have made unlearning an increasingly active research direction.
@@ 2 Related Works
Several approaches to machine unlearning have been explored. A straightforward strategy applies gradient ascent on a designated forget-set to counteract learned influence. Other methods operate in latent space, steering samples toward alternative semantic directions. Weight-saliency-based approaches attempt to identify and suppress parameter subsets most responsible for memorisation. More recently, works such as Ermon et al. (SISS framework) propose relaxing the goal of exact unlearning by instead attenuating the contribution of forgotten data in a controlled manner, acknowledging that strict removal often harms generative fidelity.
@@ 4.1 Variational Autoencoders (VAEs)
$\mathcal{L}_{\text{VAE}}(\theta,\phi;\mathbf{x})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\Big[\log p_{\theta}(\mathbf{x}|\mathbf{z})\Big]-D_{\text{KL}}\Big(q_{\phi}(\mathbf{z}|\mathbf{x})\parallel p(\mathbf{z})\Big)$
@@ 4.2.1 DDPM (VAE-Based Approach)
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}\Big(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\Big)$
@@
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}\Big(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)\Big)$
@@
$\mathcal{L}_{\text{DDPM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\big(\sqrt{\bar{\alpha}_{t}}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon},t\big)\right\|_{2}^{2}\right]$
@@ 4.2.2 Score-Based DDPM & SDE Approach
$\mathcal{L}_{\text{SGM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t}}\left[\left\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}\mid\mathbf{x}_{0})\right\|_{2}^{2}\right]$
@@ 4.3 SISS Framework
The SISS framework utilizes a mixture distribution $q_{\lambda}(m_{t})$ from which the noisy sample $m_{t}$ is drawn at time step $t$
$\forall\lambda\in[0,1]:q_{\lambda}(m_{t}|x,a)=(1-\lambda)q(m_{t}|x)+\lambda q(m_{t}|a)$ (2)
where $q(m_{t}|x)=\mathcal{N}(m_{t}|\gamma_{t}x,\sigma_{t}^{2}I)$. It is like tossing a coin and if it was heads which happens with probability of $\lambda$, we sample from $\mathcal{F}$, otherwise, we sample from $\mathcal{R}=\mathcal{D}\backslash\mathcal{F}$.
@@
$\mathcal{L}_{\text{retain}}(x,a,m_{t},t;\theta)$ $=\frac{n}{n-k}\frac{q(m_{t}|x)}{q_{\lambda}(m_{t}|x,a)}\left\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$ (3)
$\mathcal{L}_{\text{forget}}(x,a,m_{t},t;\theta)$ $=\frac{k}{n-k}\frac{q(m_{t}|a)}{q_{\lambda}(m_{t}|x,a)}\left\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$
@@
$\mathcal{L}_{\text{siss}}$ $=\mathbb{E}_{x\sim P_{X}}\,\mathbb{E}_{a\sim P_{A}}\,\mathbb{E}_{t\sim P_{T}}\,\mathbb{E}_{m_{t}\sim q_{\lambda}}\Big[\mathcal{L}_{\text{retain}}-(1+s)\,\mathcal{L}_{\text{forget}}\Big]$ (4)
$=\mathbb{E}_{x\sim P_{X}}\,\mathbb{E}_{a\sim P_{A}}\,\mathbb{E}_{t\sim P_{T}}\,\mathbb{E}_{m_{t}\sim q_{\lambda}}\Bigg[\frac{n}{n-k}\cdot\frac{q(m_{t}\mid x)}{q_{\lambda}(m_{t}\mid x,a)}\Big\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\Big\|_{2}^{2}$
$\quad-(1+s)\cdot\frac{k}{n-k}\cdot\frac{q(m_{t}\mid a)}{q_{\lambda}(m_{t}\mid x,a)}\Big\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\Big\|_{2}^{2}\Bigg]$
@@
Algorithm 1: Static-$\lambda$ SISS Unlearning for DDPMs  1:Dataset $\mathcal{D}$, Forget set $\mathcal{F}$, Retain set $\mathcal{R}$, Model $\varepsilon_{\theta}$, Noise $q$, mixture weight $\lambda$, learning rate $\eta$, scaling $s$, epochs $E$ 2:for $e=1$ to $E$ do 3:  Sample $x\sim\mathcal{R}$, $a\sim\mathcal{F}$, $t\sim P_{T}$ 4:  Sample latent: $m_{t}\sim q_{\lambda}(m_{t})=(1-\lambda)q(m_{t}|x)+\lambda q(m_{t}|a)$ 5:  Retain Loss: 6:   $\mathcal{L}_{\mathrm{ret}}\leftarrow\frac{|\mathcal{D}|}{|\mathcal{R}|}\frac{q(m_{t}|x)}{q_{\lambda}(m_{t})}\left\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\varepsilon_{\theta}(m_{t},t)\right\|^{2}$ 7:  Forget Loss: 8:   $\mathcal{L}_{\mathrm{forget}}\leftarrow\frac{|\mathcal{D}|}{|\mathcal{F}|}\frac{q(m_{t}|a)}{q_{\lambda}(m_{t})}\left\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\varepsilon_{\theta}(m_{t},t)\right\|^{2}$ 9:  Total loss: $\mathcal{L}_{\mathrm{SISS}}\leftarrow\mathcal{L}_{\mathrm{ret}}-(1+s)\mathcal{L}_{\mathrm{forget}}$ 10:  Optionally clip/scale gradient of $\mathcal{L}_{\mathrm{forget}}$ 11:  Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\mathrm{SISS}}$ 12:end for
@@ 5.1.1 Variational Inference for $\lambda$
$\mathcal{L}_{\text{ELBO}}$ $=\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\Big[\mathcal{L}_{\text{siss}}(\lambda)+\beta D_{\text{KL}}\Big(q_{\phi}(\lambda)||p(\lambda)\Big)\Big]$ (8)
$=\mathbb{E}_{x\sim P_{X}}\mathbb{E}_{a\sim P_{A}}\mathbb{E}_{t\sim P_{T}}\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\mathbb{E}_{m_{t}\sim q_{\lambda}}\Bigg[$
$\quad\frac{n}{n-k}\cdot\frac{q(m_{t}|x)}{q_{\lambda}(m_{t}|x,a)}\left\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$
$\quad-(1+s)\cdot\frac{k}{n-k}\cdot\frac{q(m_{t}|a)}{q_{\lambda}(m_{t}|x,a)}\left\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}\Bigg]+\beta D_{\text{KL}}\Big(q_{\phi}(\lambda)||p(\lambda)\Big)$
@@ 5.1.2 Context Feature Vector and Reparameterization
$z=\mu_{\phi}(v)+\xi\cdot\sigma_{\phi}(v),\quad\xi\sim\mathcal{N}(\vec{0},I)$
@@ 5.1.3 Backpropagation and Gradient Update
$\frac{\partial\mathcal{L}_{\text{ELBO}}}{\partial\phi}=\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\Bigg[\frac{\partial\mathcal{L}_{\text{siss}}}{\partial\lambda}\cdot\frac{\partial\lambda}{\partial z}\Bigg(\frac{\partial z}{\partial\mu}\cdot\frac{\partial\mu}{\partial\phi}+\frac{\partial z}{\partial\sigma}\cdot\frac{\partial\sigma}{\partial\phi}\Bigg)+\beta\frac{\partial D_{\text{KL}}}{\partial\phi}\Bigg]$
@@
Figure 2: Our own method is depicted in this image.

### v2

@@ Abstract
Machine Unlearning (MU) is critical for large generative models, such as VAEsand DDPMs,to enforce the right to be forgotten and ensure safe content generation without expensive retraining. Existing methods, like the Static-$\lambda$ SISS frameworkfor Diffusion Models, rely on a fixed data mixing parameter $\lambda$, which is suboptimal as the necessary unlearning rate varies dynamically across samples and training stages.
We introduce Adaltive-$\lambda$ SISS, a novel approach that transforms the mixing parameter $\lambda$ from a fixed constant into a latent variable whose optimal value is dynamically inferred at each unlearning step. We leverage a Variational Inference-based objective, employing a small neural network to learn an adaptive posterior distribution $q_{\phi}(\lambda)$. This network conditions on context features derived from the SISS loss components, enabling simultaneous and robust optimization of the diffusion model and the $\lambda$-inference mechanism.
@@ 1 Introduction
Practical motivations for unlearning span several dimensions. Individuals may demand that facial images resembling them in datasets such as CelebA be removed from the learned model. More broadly, erasing semantic concepts or class labels enables a pathway toward controllable editing [24] and concept-level modification within generative models. In addition, unlearning supports safe guidance, where a model avoids sampling from inappropriate or undesirable regions of the data distribution. These considerations, together with the infeasibility of retraining large models from scratch, have made unlearning an increasingly active research direction.
@@ 2 Related Works
Several approaches to machine unlearning have been explored. A straightforward strategy applies gradient ascent[21] on a designated forget-set to counteract learned influence. Other methods operate in latent space, steering samples toward alternative semantic directions. Weight-saliency-based approaches attempt to identify and suppress parameter subsets most responsible for memorization. More recently, works such as Ermon et al. (SISS framework) propose relaxing the goal of exact unlearning by instead attenuating the contribution of forgotten data in a controlled manner, acknowledging that strict removal often harms generative fidelity.
@@ 4.1 Variational Autoencoders (VAEs)
$\mathcal{L}_{\text{VAE}}(\theta,\phi;\mathbf{x})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\Big[\log p_{\theta}(\mathbf{x}|\mathbf{z})\Big]-D_{\text{KL}}\Big(q_{\phi}(\mathbf{z}|\mathbf{x})\parallel p(\mathbf{z})\Big)$ (1)
@@ 4.2.1 DDPM (VAE-Based Approach)
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}\Big(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\Big)$ (2)
@@
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}\Big(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)\Big)$ (3)
@@
$\mathcal{L}_{\text{DDPM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\big(\sqrt{\bar{\alpha}_{t}}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon},t\big)\right\|_{2}^{2}\right]$ (4)
@@ 4.2.2 Score-Based DDPM & SDE Approach
$\mathcal{L}_{\text{SGM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t}}\left[\left\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}\mid\mathbf{x}_{0})\right\|_{2}^{2}\right]$ (5)
@@ 4.3 SISS Framework
The SISS framework utilizes a mixture distribution $q_{\lambda}(m_{t}|x_{r},x_{f})$ from which the noisy sample $m_{t}$ is drawn at time step $t$ (assume that $x_{r}$ is from retain-set and $x_{f}$ is from forget-set)
$\forall\lambda\in[0,1]:q_{\lambda}(m_{t}|x_{r},x_{f})=(1-\lambda)q(m_{t}|x_{r})+\lambda q(m_{t}|x_{f})$ (7)
where $q(m_{t}|x)=\mathcal{N}(m_{t}|\gamma_{t}x,\sigma_{t}^{2}I)$. It is like tossing a coin and if it was heads which happens with probability of $\lambda$, we sample from $\mathcal{F}$, otherwise, we sample $m_{t}$ from $\mathcal{R}=\mathcal{D}\backslash\mathcal{F}$.
@@
$\mathcal{L}_{\text{retain}}(x_{r},x_{f},m_{t},t;\theta)$ $=\frac{n}{n-k}\frac{q(m_{t}|x)}{q_{\lambda}(m_{t}|x_{r},x_{f})}\left\|\frac{m_{t}-\gamma_{t}x_{r}}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$ (8)
$\mathcal{L}_{\text{forget}}(x_{r},x_{f},m_{t},t;\theta)$ $=\frac{k}{n-k}\frac{q(m_{t}|a)}{q_{\lambda}(m_{t}|x_{r},x_{f})}\left\|\frac{m_{t}-\gamma_{t}x_{f}}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$
@@
$\mathcal{L}_{\text{siss}}$ $=\mathbb{E}_{x_{r}\sim P_{R}}\,\mathbb{E}_{x_{f}\sim P_{F}}\,\mathbb{E}_{t\sim P_{T}}\,\mathbb{E}_{m_{t}\sim q_{\lambda}}\Big[\mathcal{L}_{\text{retain}}-(1+s)\,\mathcal{L}_{\text{forget}}\Big]$ (9)
$=\mathbb{E}_{x_{r}\sim P_{R}}\,\mathbb{E}_{x_{f}\sim P_{F}}\,\mathbb{E}_{t\sim P_{T}}\,\mathbb{E}_{m_{t}\sim q_{\lambda}}\Bigg[\frac{n}{n-k}\cdot\frac{q(m_{t}\mid x_{r})}{q_{\lambda}(m_{t}\mid x_{r},x_{f})}\Big\|\frac{m_{t}-\gamma_{t}x_{r}}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\Big\|_{2}^{2}$
$\quad-(1+s)\cdot\frac{k}{n-k}\cdot\frac{q(m_{t}\mid x_{f})}{q_{\lambda}(m_{t}\mid x_{r},x_{f})}\Big\|\frac{m_{t}-\gamma_{t}x_{f}}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\Big\|_{2}^{2}\Bigg]$
@@
Algorithm 1: Static-$\lambda$ SISS Unlearning for DDPMs  1:Dataset $\mathcal{D}$, Forget set $\mathcal{F}$, Retain set $\mathcal{R}$, Model $\varepsilon_{\theta}$, Noise $q$, mixture weight $\lambda$, learning rate $\eta$, scaling $s$, epochs $E$ 2:for $e=1$ to $E$ do 3:  Sample $x_{r}\sim\mathcal{R}$, $x_{f}\sim\mathcal{F}$, $t\sim P_{T}$ 4:  Sample latent: $m_{t}\sim q_{\lambda}(m_{t}\mid x_{r},x_{f})=(1-\lambda)q(m_{t}|x_{r})+\lambda q(m_{t}\mid x_{f})$ 5:  Retain Loss: 6:   $\mathcal{L}_{\mathrm{retain}}\leftarrow\frac{|\mathcal{D}|}{|\mathcal{R}|}\frac{q(m_{t}|x_{r})}{q_{\lambda}(m_{t}|x_{r},x_{f})}\left\|\frac{m_{t}-\gamma_{t}x_{r}}{\sigma_{t}}-\varepsilon_{\theta}(m_{t},t)\right\|^{2}$ 7:  Forget Loss: 8:   $\mathcal{L}_{\mathrm{forget}}\leftarrow\frac{|\mathcal{D}|}{|\mathcal{F}|}\frac{q(m_{t}|x_{f})}{q_{\lambda}(m_{t}|x_{r},x_{f})}\left\|\frac{m_{t}-\gamma_{t}x_{f}}{\sigma_{t}}-\varepsilon_{\theta}(m_{t},t)\right\|^{2}$ 9:  Total loss: $\mathcal{L}_{\mathrm{SISS}}\leftarrow\mathcal{L}_{\mathrm{retain}}-(1+s)\mathcal{L}_{\mathrm{forget}}$ 10:  Optionally clip/scale gradient of $\mathcal{L}_{\mathrm{forget}}$ 11:  Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\mathrm{SISS}}$ 12:end for
@@ 5.1.1 Variational Inference for $\lambda$
$\mathcal{L}_{\text{ELBO}}$ $=\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\Big[\mathcal{L}_{\text{siss}}(\lambda)\Big]+\beta D_{\text{KL}}\Big(q_{\phi}(\lambda)||p(\lambda)\Big)$ (13)
$=\mathbb{E}_{x_{r}\sim P_{R}}\mathbb{E}_{x_{f}\sim P_{F}}\mathbb{E}_{t\sim P_{T}}\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\mathbb{E}_{m_{t}\sim q_{\lambda}}\Bigg[$
$\quad\frac{n}{n-k}\cdot\frac{q(m_{t}|x_{r})}{q_{\lambda}(m_{t}|x_{r},x_{f})}\left\|\frac{m_{t}-\gamma_{t}x_{r}}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$
$\quad-(1+s)\cdot\frac{k}{n-k}\cdot\frac{q(m_{t}|x_{f})}{q_{\lambda}(m_{t}|x_{r},x_{f})}\left\|\frac{m_{t}-\gamma_{t}x_{f}}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}\Bigg]+\beta D_{\text{KL}}\Big(q_{\phi}(\lambda)||p(\lambda)\Big)$
@@ 5.1.2 Context Feature Vector and Reparameterization
$z=\mu_{\phi}(v)+\xi\cdot\sigma_{\phi}(v),\quad\xi\sim\mathcal{N}(0,1)$
@@ 5.1.3 Backpropagation and Gradient Update
$\frac{\partial\mathcal{L}_{\text{ELBO}}}{\partial\phi}=\mathbb{E}_{\xi\sim\mathcal{N}(0,1)}\Bigg[\frac{\partial\mathcal{L}_{\text{siss}}}{\partial\lambda}\cdot\frac{\partial\lambda}{\partial z}\Bigg(\frac{\partial z}{\partial\mu}\cdot\frac{\partial\mu}{\partial\phi}+\frac{\partial z}{\partial\sigma}\cdot\frac{\partial\sigma}{\partial\phi}\Bigg)+\beta\frac{\partial D_{\text{KL}}}{\partial\phi}\Bigg]$ (14)
@@
Figure 2: In this image the process that how our model works is depited. the model uses the retain and forget losses of SISS and uses them to encode a latent variable $z$, which later by sampling from posterior distribution we get $z$ and by feeding it into a sigmoid function we get the $\lambda$ we wanted.